OpenStudy (anonymous):

Solve the foloowing proportion for x: 2/11 = X/91-X a.14 b.20.22 c.70.78 d.77

OpenStudy (angela210793):

2(91-x)=11x 182-2x=11x 182=13x x=182/13

OpenStudy (anonymous):

11x=182-2x 13x=182 x=14 ans a